# The Timescales of Embedded Star Formation as Observed in STARFORGE

### v1

@@ 6.1.3 Interpreting Measures of Embeddedness for Semi/Un-resolved Association and Star Forming Regions
We find that the overall timescales for formation and emergence scale strongly with the initial free fall time, such that the larger mass cloud has a larger characteristic timescales. We note however, that the simulations analyzed here only consider two masses with extremely similar initial conditions, such that the main difference is cloud mass and density. We also note that these timescales reflected the association as a whole, and that one might well derive different timescales when analyzing distinct sub-regions. In fact, when considering the embeddedness of individual stars in both clouds, we see the behavior of smaller stars staying embedded longer than massive stars (Section 4.3. More instantaneous formation of stars would then manifest in the more massive complexes (i.e., higher number of massive stars) having shorter embedded timescales.
6.2 Model Limitations
@@ 7 Conclusion
This work was completed as a part of the pre-doctoral fellowship at the Center for Computational Astrophysics. TW thanks the Flatiron Institute and the Simons Foundation. The Flatiron Institute is funded by the Simons Foundation. We graciously thank Angela Adamo, Cara Battersby, Lise Ramambason and Philip Myers for thoughtful conversations which helped shape this work.

### v2

@@ VI.1.3 Interpreting Measures of Embeddedness for Semi/Un-resolved Association and Star Forming Regions
We find that the overall timescales for formation and emergence scale strongly with the initial free fall time, such that the larger mass cloud has a larger characteristic timescales. We note however, that the simulations analyzed here only consider two masses with extremely similar initial conditions, such that the main difference is cloud mass and density. We also note that these timescales reflected the association as a whole, and that one might well derive different timescales when analyzing distinct sub-regions. In fact, when considering the embeddedness of individual stars in both clouds, we see the behavior of smaller stars staying embedded longer than massive stars (Section IV.3). More instantaneous formation of stars would then manifest in the more massive complexes (i.e., higher number of massive stars) having shorter embedded timescales.
VI.2 Model Limitations
@@
Similarly with regards to the galactic context, the simulations do not include a background galactic potential or vertical stratification, and therefore lack the additional confining weight and external pressure present in real galactic discs. In vertically stratified disc simulations, feedback-driven bubbles evolve in balance with the vertical weight of the ISM, and their ability to clear gas locally depends on this balance (e.g., Faucher-Giguère et al., 2013; Ostriker & Kim, 2022). As such, stronger vertical gravity or higher ambient pressure can prolong the time that gas remains bound and dense, potentially lengthening the embedded phase. Our results should be viewed as characterizing the locally driven de-embedding physics, with the understanding that large-scale galactic context can modulate the absolute timescales.
@@ VII Conclusion
Both 24µm, and H emission are brightest during the short interval when dense gas and dust remain tightly coupled to massive stars. At later times, dust and H luminosities decline even as the ionized gas mass grows, enforcing the local clearing produced by high-mass stars as the dominant mechanism in setting embedded timescales. Additionally, the duration of the 24 µmdominance agrees well with observationally inferred embedded timescales, reinforcing that the simulations correspond to the same physics that governs embedded tracers in real systems.